A merchant mixed 10 lb of a cinnamon tea with 2 lb of spice tea. The 12-pound mixture cost $38. A second mixture included 16 lb of the cinnamon tea and 8 lb of the spice tea. The 24-pound mixture cost $80. Find the cost per pound of the cinnamon tea and of the spice tea.
Answer by Boreal(15235)

10x+2y=38, where x is the dollar cost of cinnamon and y the dollar cost of the spice.
16x+8y=80
multiply the top by -4
-40x-8y=-152
16x+8y=80. Add them and y is eliminated.
-24x=-72
x=$3
go to first equation and 30+2y=38, substituting $3 for x.
2y=8
y=$4.
Check using second equation.
48+32=80
Cinnamon is $3 per pound
Spice is $4 per pound.